whenever horses a,b and c race together , their respective probabilities of winning are 0.3, 0.5 and 0.2. they race three times .

(a) Find the probability that the same horse wins all three races .

(b) Find the probability that a,b,,c each win one race.


1
Expert's answer
2022-02-15T14:11:17-0500

Given "P(a)=0.3, P(b)=0.5, P(c)=0.2"

(a)

"P(aaa)=(P(a))^3=(0.3)^3=0.027"

"P(bbb)=(P(b))^3=(0.5)^3=0.125"

"P(ccc)=(P(c))^3=(0.2)^3=0.008"

"P(one\ wins\ all)=0.027+0.125+0.008=0.16"

(b)

First race "a", second race "b", third race "c"

"P(abc)=0.3(0.5)(0.2)=0.03"

First race "a", second race "c", third race "b"

"P(abc)=0.3(0.2)(0.5)=0.03"

First race "b", second race "a", third race "c"

"P(bac)=0.5(0.3)(0.2)=0.03"

First race "b", second race "c", third race "a"

"P(bca)=0.5(0.2)(0.3)=0.03"

First race "c", second race "a", third race "b"

"P(cab)=0.2(0.3)(0.5)=0.03"

First race "c", second race "b", third race "a"

"P(cba)=0.2(0.5)(0.3)=0.03"




"P(each\ wins\ one\ race)=6(0.03)=0.18"